OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (dtan5457):

f(g(x)) means to put g's value into f. f=9x-2 g=-x+3 f(g(x))=9(-x+3)-2 Simplify it now